# Deep Sparse Latent Feature Models for Knowledge Graph Completion

### v1

@@ Abstract
Recent progress in knowledge graph completion (KGC) has focused on text-based approaches to address the challenges of large-scale knowledge graphs (KGs). Despite their achievements, these methods often overlook the intricate interconnections between entities, a key aspect of the underlying topological structure of a KG. Stochastic blockmodels (SBMs), particularly the latent feature relational model (LFRM), offer robust probabilistic frameworks that can dynamically capture latent community structures and enhance link prediction. In this paper, we introduce a novel framework of sparse latent feature models for KGC, optimized through a deep variational autoencoder (VAE). Our approach not only effectively completes missing triples but also provides clear interpretability of the latent structures, leveraging textual information. Comprehensive experiments on the WN18RR, FB15k-237, and Wikidata5M datasets show that our method significantly improves performance by revealing latent communities and producing interpretable representations.
@@ 1 Introduction
The majority of real-world phenomena exhibit multifaceted characteristics. For instance, social networks are not merely a collection of isolated individuals but represent a complex web of interactions across various contexts. Knowledge graphs (KGs) organize information into triples $(h,r,t)$, where $h$ denotes the head entity, $t$ the tail entity, and $r$ the relationship, forming extensive semantic networks. However, real-world KGs like DBpedia (Auer et al., 2007) and Wikidata (Vrandečić & Krötzsch, 2014) often suffer from incompleteness, missing key entities and relationships (Dong et al., 2014). Knowledge graph completion (KGC) aims to infer this missing information, improving the utility and completeness of the graph.
Early developments in KGC centered around knowledge graph embedding (KGE) techniques (Bordes et al., 2013; Sun et al., 2019; Balažević et al., 2019), which focused on learning low-dimensional embeddings for entities and relations, applying various scoring functions to triples. More recently, text-based methods (Yao et al., 2019; Wang et al., 2021a; 2022) utilizing pre-trained language models (PLMs) have achieved state-of-the-art performance on large-scale datasets such as Wikidata5M (Wang et al., 2021b). These approaches generally rely on the transformation of the head embedding $\mathbf{h}$ into the tail $\mathbf{t}$ through the relation $\mathbf{r}$, yet the complex interconnectivity among communities associated with entities remains insufficiently exploited. As highlighted by Stanley et al. (2019), network topologies typically exhibit dense connections within groups and fewer connections between them. Inspired by this, we approach triple completion from a broader perspective by focusing on relational connections across entity communities. For instance, in the KG depicted in Figure 1, entities are categorized into overlapping communities. To answer the query regarding the relationship between Michael Jordan and Gregg Popovich, ’acquaintance’ emerges as a plausible candidate, which can be inferred from the interconnections observed between the two communities—’NBA Players’ and ’NBA Coaches’.
@@
Uncovering the latent structure of graph data is a key area of focus in statistical network analysis (Porter et al., 2009; Latouche et al., 2010). Stochastic blockmodels (SBMs) (Airoldi et al., 2008; Miller et al., 2009; Latouche et al., 2011) are a widely recognized class of probabilistic models that assign cluster memberships to graph nodes, and are highly regarded in both academic and industrial settings. A notable variant is the latent feature relational model (LFRM), a type of overlapping stochastic blockmodel (OSBM), which allows nodes to belong to multiple groups and leverages an Indian Buffet Process (IBP) prior on the node-community assignment matrix $Z$ to discover the number of latent communities. These models typically rely on MCMC (Miller et al., 2009) or variational inference (Zhu et al., 2016) to infer latent variables. While DGLFRM (Mehta et al., 2019) enhances SBM inference using a deep sparse variational autoencoder (VAE) (Kingma & Welling, 2013), it is not tailored for KGC tasks and faces challenges when scaled to large graphs with hundreds of thousands or even millions of nodes.
@@
Contributions. We propose DSLFM-KGC, a novel method for tackling the KGC challenge by utilizing latent community structures in KGs. Our main contributions are as follows: i) we design an end-to-end probabilistic model for KGC that integrates additional sparse clustering features into triple representation, implemented through a deep VAE (Kingma & Welling, 2013); ii) DSLFM-KGC provides robust performance and interpretability in completing missing triples by leveraging community-level interconnections in entities; and iii) the deep architecture allows for scalable inference. Through extensive experiments on the WN18RR, FB15k-237, and Wikidata5M datasets, iv) we showcase our model’s superior capability and scalability in managing KGC tasks and uncovering interpretable latent structures.
@@ 2.1 Latent feature relational model
The SBMs (Holland et al., 1983; Airoldi et al., 2008; Miller et al., 2009) are fundamental approaches for analyzing relational data, where a graph with $N$ nodes is represented by a binary adjacency matrix $A\in\{0,1\}^{N\times N}$. In this matrix, $A_{i,j}=1$ indicates a link between node $i$ and node $j$. Each node $i$ in an SBM is associated with a one-hot latent variable $\mathbf{z}_{i}\in\{0,1\}^{K}$ to indicate its community membership, where $K$ is the number of node communities.
For scenarios where nodes belong to multiple communities, the OSBM (Latouche et al., 2011) adapts the latent indicator $\mathbf{z}_{i}$ into a multivariate Bernoulli vector consisting of $K$ independent Bernoulli variables, denoted as $\mathbf{z}_{i}\sim\mathcal{MB}(\mathbf{z}|\bm{\pi})$:
@@
Expanding on OSBM, the LFRM integrates the IBP prior (Miller et al., 2009) on the binary node-community matrix $Z=[\mathbf{z}_{1},\ldots,\mathbf{z}_{N}]^{\top}$, enabling dynamic learning of the number of communities. Traditional inference methods used in SBMs, such as MCMC or variational inference, often struggle to scale in large networks. To overcome this, DGLFRM (Mehta et al., 2019) uses a VAE (Kingma & Welling, 2013), employing a graph convolutional network (GCN) (Kipf & Welling, 2016) to encode the variational distribution $q(Z)$ and a non-linear multilayer perceptron (MLP) to model the link probability $p(A_{i,j}|\mathbf{z}_{i},\mathbf{z}_{j},W)$. Despite its advances, DGLFRM encounters difficulties when applied to large-scale heterogeneous KGs, which feature entities and relations of diverse types.
@@ 2.2 Knowledge graph completion
A KG is commonly defined as $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T})$, where $\mathcal{E}$ is the set of entities, and $\mathcal{R}$ is the set of relations. The set $\mathcal{T}=\{(h,r,t)|h,t\in\mathcal{E},r\in\mathcal{R}\}$ contains factual triples, each representing a directed labeled edge $h\mathop{\rightarrow}\limits^{r}t$ in the KG. Furthermore, modern KGs often include meta-information $\mathcal{M}$, such as natural language descriptions (Yao et al., 2019; Wang et al., 2022) or multi-modal data (Zhang et al., 2024a). For any entity $e\in\mathcal{E}$ and any relation $r\in\mathcal{R}$, $\mathcal{M}(e)$ and $\mathcal{M}(r)$ denote the corresponding meta-information.
For a given query $(h,r,?)$, the task of KGC involves identifying the missing tail entity by retrieving the most plausible candidate $\hat{t}$ from the entity set $\mathcal{E}$, such that $(h,r,\hat{t})$ is valid. From the KGC perspective, we model a KG as containing a query set $\mathcal{Q}=\{(h,r)|h\in\mathcal{E},r\in\mathcal{R}\}$, a candidate answer (entity) set $\mathcal{E}$, and a mapping $\mathcal{A}:\mathcal{Q}\times\mathcal{E}\rightarrow\{0,1\}$ that identifies whether a query has a valid answer in the KG $\mathcal{G}$. This mapping is represented as a binary matrix $A\in\{0,1\}^{|\mathcal{Q}|\times|\mathcal{E}|}$, analogous to an adjacency matrix, where $A_{hr,t}=1$ if the triple $(h,r,t)\in\mathcal{T}$, and $A_{hr,t}=0$ otherwise.
@@ 3 Methodology
This section presents the framework of DSLFM-KGC. We begin by describing the probabilistic generative model for KGs, emphasizing its application to KGC. Following this, we elaborate the VAE architecture employed for inference, detailing the design and implementation of both the encoder and decoder.
@@ 3.1 Generative model
We assume that triples within a KG are conditionally independent, given their latent communities. The generation process of a KG unfolds as follows:
@@
Let $Z_{\text{qry}}$ and $Z_{\text{ans}}$ denote the membership indicator matrices for queries and answers, respectively, and let $W_{\text{qry}}$ and $W_{\text{ans}}$ denote the membership strength matrices. Then, $F_{\text{qry}}=Z_{\text{qry}}\odot W_{\text{qry}}$ and $F_{\text{ans}}=Z_{\text{ans}}\odot W_{\text{ans}}$ constitute a sparse latent feature model (Ghahramani & Griffiths, 2005; d’Aspremont et al., 2004; Jolliffe et al., 2003). We use the Indian Buffet Process (IBP) (Griffiths & Ghahramani, 2011) prior on the indicator matrices to facilitate the learning of the number of communities, thereby establishing an infinite latent feature model (Ghahramani & Griffiths, 2005).
@@ 3.2 VAE encoder
We adopt the stick-breaking construction of the IBP (Teh et al., 2007) to model $\mathbf{z}_{hr}$:
@@
$q_{\phi}(\mathcal{H})=\prod_{k=1}^{K}\left(\prod_{(h,r)\in\mathcal{Q}}q_{\phi}(v_{hr,k})q_{\phi}(z_{hr,k})q_{\phi}(w_{hr,k})\prod_{t\in\mathcal{A}}q_{\phi}(v_{t,k})q_{\phi}(z_{t,k})q_{\phi}(w_{t,k})\right)$ (10)
@@
$q_{\phi}(v_{hr,k})$ $\triangleq\text{Beta}(c_{hr,k},d_{hr,k}),~{}~{}k=1,\ldots,K$ (11)
$q_{\phi}(z_{hr,k})$ $\triangleq q_{\phi}(z_{hr,k}|\mathcal{Q})=\text{Bernoulli}(\pi_{hr,k}(\mathcal{Q})),~{}~{}k=1,\ldots,K$ (12)
$q_{\phi}(w_{hr,k})$ $\triangleq q_{\phi}(w_{hr,k}|\mathcal{Q})=\mathcal{N}(\mu_{hr,k}(\mathcal{Q}),\sigma^{2}_{hr,k}(\mathcal{Q})),~{}~{}k=1,\ldots,K$ (13)
where $\pi_{hr,k},\mu_{hr,k}$ and $\sigma^{2}_{hr,k}$ are outputs of the encoder network $h_{\phi}$, i.e., $\{\bm{\pi}_{hr},\bm{\mu}_{hr},\bm{\sigma}^{2}_{hr}\}_{(h,r)\in\mathcal{Q}}=h_{\phi}(\mathcal{Q})$ with $\mathcal{Q}$ as the input. In experiment, we found that treating $\mathbf{c}_{hr}$ and $\mathbf{d}_{hr}$ as part of the encoder parameters (instead of encoding them from the posterior) helps mitigate over-parameterization. We define $q_{\phi}(v_{t,k})$, $q_{\phi}(z_{t,k})$ and $q_{\phi}(w_{t,k})$ in a similar vein, with $\mathcal{Q}$ replaced by $\mathcal{E}$.
Following recent progress in text-based approaches for the KGC task (Yao et al., 2019; Wang et al., 2022), we employ the strategy that individually encodes the textual descriptions of queries and answers using two BERT (Devlin et al., 2019) encoders, sharing pre-trained weights, and applying mean pooling:
$\mathbf{e}_{hr}=\text{Pool}(\text{BERT}_{\text{qry}}(\mathbf{x}_{hr})),\quad\mathbf{e}_{t}=\text{Pool}(\text{BERT}_{\text{ans}}(\mathbf{x}_{t}))$ (14)
@@
$\{\pi_{hr,k},\mu_{hr,k},\sigma_{hr,k}\}_{k=1}^{K}=\text{MLP}(\mathbf{e}_{hr}),\quad\{\pi_{t,k},\mu_{t,k},\sigma_{t,k}\}_{k=1}^{K}=\text{MLP}(\mathbf{e}_{t})$ (15)
It is important to note that our approach leverages latent structures encoded within the textual semantic space, which has demonstrated enhanced expressiveness for KGC tasks. Additionally, the flexibility of our model allows for the use of various types of encoders, such as a multi-modal one, to further enhance expressiveness. We plan to explore these possibilities in future research.
We denote the overall encoder network with parameters $\phi$ as $h_{\phi}$. Integrating textual inputs not only enhances the performance of our model, but also provides deeper insights into the latent structure. This allows for the exploration of the mined communities through the descriptions of the entities within, the benefits of which will be demonstrated in the experiment section.
Figure 2: An overview of our DSLFM-KGC framework during inference. Initially, the encoder network $h_{\phi}$ encodes the textual information of a triple ($\mathbf{x}_{hr}$ and $\mathbf{x}_{t}$) into posterior distributions, as defined in Equations 14 and 15. Latent variables (e.g., $\mathbf{z}_{hr}$ and $\mathbf{w}_{hr}$) are then sampled using reparameterization tricks (see Appendix A.4), after which the decoder $g_{\theta}$ generates representations for the query and answer ($\mathbf{g}_{hr}$ and $\mathbf{g}_{t}$).

### v2

@@ Abstract
Recent advances in knowledge graph completion (KGC) have emphasized text-based approaches to navigate the inherent complexities of large-scale knowledge graphs (KGs). While these methods have achieved notable progress, they frequently struggle to fully incorporate the global structural properties of the graph. Stochastic blockmodels (SBMs), especially the latent feature relational model (LFRM), offer robust probabilistic frameworks for identifying latent community structures and improving link prediction. This paper presents a novel probabilistic KGC framework utilizing sparse latent feature models, optimized via a deep variational autoencoder (VAE). Our proposed method dynamically integrates global clustering information with local textual features to effectively complete missing triples, while also providing enhanced interpretability of the underlying latent structures. Extensive experiments on four benchmark datasets with varying scales demonstrate the significant performance gains achieved by our method.
@@ 1 Introduction
The majority of real-world phenomena exhibit multifaceted characteristics. For instance, social networks are not merely a collection of isolated individuals but represent a complex web of interactions across various contexts. Knowledge graphs (KGs) organize information into triples $(h,r,t)$, where $h$ denotes the head entity, $t$ the tail entity, and $r$ the relationship, forming extensive semantic networks. However, real-world KGs like DBpedia (auer2007dbpedia, ) and Wikidata (vrandevcic2014wikidata, ) often suffer from incompleteness, missing key entities and relationships (dong2014knowledge, ). Knowledge graph completion (KGC) aims to infer this missing information, improving the utility and completeness of the graph.
Early KGC research focused on knowledge graph embedding (KGE) techniques (bordes2013translating, ; sun2018rotate, ; balavzevic2019tucker, ), which aimed to learn low-dimensional embeddings for entities and relations, employing various scoring functions to evaluate triples. More recently, text-based methods leveraging pre-trained language models (PLMs) (yao2019kg, ; wang2021structure, ; wang2022simkgc, ) have attained state-of-the-art results on large-scale datasets such as Wikidata5M (wang2021kepler, ). These approaches typically rely on transforming the head embedding $\mathbf{h}$ into the tail $\mathbf{t}$ through the relation $\mathbf{r}$. However, textual descriptions often suffer from information scarcity due to length constraints. Consequently, global information, particularly the complex interrelationships among entity communities, remains underexploited. As noted by stanley2019stochastic , network topologies frequently exhibit dense intra-group connections and sparser inter-group connections. While graph neural networks (GNNs), particularly message passing neural networks (MPNNs), enhance graph representation learning by incorporating node neighborhood structures, they have been shown to underutilize neighborhood information in later studies (zhang2022rethinking, ; li2023message, ). Moreover, GNN-based KGC methods also fail to incorporate inherent clustering properties. Motivated by these observations, we propose a triple completion method that integrates global clustering information with local textual features.
@@
Uncovering the latent structure of graph data is a central focus in statistical network analysis (porter2009communities, ; latouche2010bayesian, ). Stochastic blockmodels (SBMs) (airoldi2008mixed, ; miller2009nonparametric, ; latouche2011overlapping, ) constitute a widely recognized class of probabilistic models that assign cluster memberships to graph nodes and are highly regarded in both academic and industrial contexts. A prominent variant is the latent feature relational model (LFRM), which permits nodes to belong to multiple groups and employs an Indian Buffet Process (IBP) prior on the node-community assignment matrix $Z$ to determine the number of latent communities. These models typically employ Markov chain Monte Carlo (MCMC) (miller2009nonparametric, ) or variational inference (zhu2016max, ) for latent variable inference. While DGLFRM (mehta2019stochastic, ) enhances SBM inference by utilizing a deep sparse variational autoencoder (VAE) (kingma2013auto, ), it is not specifically designed for KGC tasks and encounters difficulties when scaling to large graphs with hundreds of thousands or even millions of nodes.
@@
Contributions. We propose DSLFM-KGC, a novel method for addressing the KGC challenge by utilizing latent community structures in KGs. Our main contributions are as follows: i) we design an end-to-end probabilistic model for KGC that integrates supplementary sparse clustering features into triple representation, implemented through a deep VAE (kingma2013auto, ); ii) DSLFM-KGC exhibits robust performance and interpretability in completing missing triples by leveraging community-level interconnections in entities; and iii) the deep architecture facilitates scalable inference. Through extensive experiments on the UMLS, WN18RR, FB15k-237, and Wikidata5M datasets, iv) we showcase our model’s superior capability and scalability in managing KGC tasks and uncovering interpretable latent structures.
@@ 2.1 Latent feature relational model
The SBMs (holland1983stochastic, ; airoldi2008mixed, ; miller2009nonparametric, ) are fundamental approaches for analyzing relational data, where a graph with $N$ nodes is represented by a binary adjacency matrix $A\in\{0,1\}^{N\times N}$. In this matrix, $A_{i,j}=1$ indicates a link between node $i$ and node $j$. Each node $i$ in an SBM is associated with a one-hot latent variable $\mathbf{z}_{i}\in\{0,1\}^{K}$ to indicate its community membership, where $K$ is the number of node communities.
For scenarios where nodes belong to multiple communities, the OSBM (latouche2011overlapping, ) adapts the latent indicator $\mathbf{z}_{i}$ into a multivariate Bernoulli vector consisting of $K$ independent Bernoulli variables, denoted as $\mathbf{z}_{i}\sim\mathcal{MB}(\mathbf{z}|\bm{\pi})$:
@@
Expanding on OSBM, the LFRM integrates the IBP prior (miller2009nonparametric, ) on the binary node-community matrix $Z=[\mathbf{z}_{1},\ldots,\mathbf{z}_{N}]^{\top}$, enabling dynamic learning of the number of communities. Traditional inference methods used in SBMs, such as MCMC or variational inference, often struggle to scale in large networks. To overcome this, DGLFRM (mehta2019stochastic, ) uses a VAE (kingma2013auto, ), employing a graph convolutional network (GCN) (kipf2016semi, ) to encode the variational distribution $q(Z)$ and a non-linear multilayer perceptron (MLP) to model the link probability $p(A_{i,j}|\mathbf{z}_{i},\mathbf{z}_{j},W)$. Despite its advances, DGLFRM encounters difficulties when applied to large-scale heterogeneous KGs, which feature entities and relations of diverse types.
@@ 2.2 Knowledge graph completion
A KG is commonly defined as $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{T})$, where $\mathcal{E}$ is the set of entities, and $\mathcal{R}$ is the set of relations. The set $\mathcal{T}=\{(h,r,t)|h,t\in\mathcal{E},r\in\mathcal{R}\}$ contains factual triples, each representing a directed labeled edge $h\mathop{\rightarrow}\limits^{r}t$ in the KG. Moreover, modern KGs often include meta-information $\mathcal{M}$, such as natural language descriptions (yao2019kg, ; wang2022simkgc, ) or multi-modal data (zhang2024native, ). For any entity $e\in\mathcal{E}$ and any relation $r\in\mathcal{R}$, $\mathcal{M}(e)$ and $\mathcal{M}(r)$ denote the corresponding meta-information.
For a given query $(h,r,?)$, the task of KGC entails identifying the missing tail entity by retrieving the most plausible candidate $\hat{t}$ from the entity set $\mathcal{E}$, such that $(h,r,\hat{t})$ is valid. From the KGC perspective, we model a KG as comprising a query set $\mathcal{Q}=\{(h,r)|h\in\mathcal{E},r\in\mathcal{R}\}$, a candidate answer (entity) set $\mathcal{E}$, and a mapping $\mathcal{A}:\mathcal{Q}\times\mathcal{E}\rightarrow\{0,1\}$ that determines whether a query has a valid answer in the KG $\mathcal{G}$. This mapping is represented as a binary matrix $A\in\{0,1\}^{|\mathcal{Q}|\times|\mathcal{E}|}$, analogous to an adjacency matrix, where $A_{hr,t}=1$ if the triple $(h,r,t)\in\mathcal{T}$, and $A_{hr,t}=0$ otherwise.
@@ 3 Methodology
This section presents the framework of DSLFM-KGC. We begin by describing the probabilistic framework for KGs, emphasizing its application to KGC. Following this, we elaborate the VAE architecture employed for inference, detailing the design and implementation of both the encoder and decoder.
@@ 3.1 Generative model
We assume that triples within a KG are conditionally independent, given their latent communities. The generative process of a KG unfolds as follows:
@@
Let $Z_{\text{qry}}$ and $Z_{\text{ans}}$ denote the membership indicator matrices for queries and answers, respectively, and let $W_{\text{qry}}$ and $W_{\text{ans}}$ denote the latent feature matrices. Then, $F_{\text{qry}}=Z_{\text{qry}}\odot W_{\text{qry}}$ and $F_{\text{ans}}=Z_{\text{ans}}\odot W_{\text{ans}}$ constitute a sparse latent feature model (ghahramani2005infinite, ; d2004direct, ; jolliffe2003modified, ). We use the Indian Buffet Process (IBP) (griffiths2011indian, ) prior on the indicator matrices to facilitate the learning of the number of communities, thereby establishing an infinite latent feature model (ghahramani2005infinite, ).
@@
Figure 1: An overview of our DSLFM-KGC framework during inference. Initially, the encoder network $h_{\phi}$ encodes the textual information of a triple ($\mathbf{x}_{hr}$ and $\mathbf{x}_{t}$) into posterior distributions, as defined in Equations 15 and 16. Latent variables (e.g., $\mathbf{z}_{hr}$ and $\mathbf{w}_{hr}$) are then sampled using reparameterization tricks (see Appendix B.4), after which the decoder $g_{\theta}$ generates representations for the query and answer ($\mathbf{g}_{hr}$ and $\mathbf{g}_{t}$).
@@ 3.2 VAE encoder
We adopt the stick-breaking construction of the IBP (teh2007stick, ) to model $\mathbf{z}_{hr}$:
@@
$q_{\phi}(\mathcal{H})=q_{\phi}(\mathcal{H}_{\text{qry}})q_{\phi}(\mathcal{H}_{\text{ans}})$ (10)
where $\mathcal{H}_{\text{qry}}$ and $\mathcal{H}_{\text{ans}}$ denote the query and answer latent variables, respectively. $q_{\phi}(\mathcal{H}_{\text{qry}})$ and $q_{\phi}(\mathcal{H}_{\text{ans}})$ are defined as follows:
$q_{\phi}(\mathcal{H}_{\text{qry}})$ $=\prod_{(h,r)\in\mathcal{Q}}\prod_{k=1}^{K}q_{\phi}(v_{hr,k})q_{\phi}(z_{hr,k})q_{\phi}(w_{hr,k})$ (11)
$q_{\phi}(\mathcal{H}_{\text{ans}})$ $=\prod_{t\in\mathcal{E}}\prod_{k=1}^{K}q_{\phi}(v_{t,k})q_{\phi}(z_{t,k})q_{\phi}(w_{t,k})$
@@
$q_{\phi}(v_{hr,k})$ $\triangleq\text{Beta}(c_{hr,k},d_{hr,k})$ (12)
$q_{\phi}(z_{hr,k})$ $\triangleq q_{\phi}(z_{hr,k}|\mathcal{Q})=\text{Bernoulli}(\pi_{hr,k}(\mathcal{Q}))$ (13)
$q_{\phi}(w_{hr,k})$ $\triangleq q_{\phi}(w_{hr,k}|\mathcal{Q})=\mathcal{N}(\mu_{hr,k}(\mathcal{Q}),\sigma^{2}_{hr,k}(\mathcal{Q}))$ (14)
where $\pi_{hr,k},\mu_{hr,k}$ and $\sigma^{2}_{hr,k}$ are outputs of the encoder network $h_{\phi}$, i.e., $\{\bm{\pi}_{hr},\bm{\mu}_{hr},\bm{\sigma}^{2}_{hr}\}_{(h,r)\in\mathcal{Q}}=h_{\phi}(\mathcal{Q})$ with $\mathcal{Q}$ as the input. In experiment, we observed that treating $\mathbf{c}_{hr}$ and $\mathbf{d}_{hr}$ as part of the encoder parameters (instead of encoding them from the posterior) helps mitigate over-parameterization. We define $q_{\phi}(v_{t,k})$, $q_{\phi}(z_{t,k})$ and $q_{\phi}(w_{t,k})$ in a similar vein, with $\mathcal{Q}$ replaced by $\mathcal{E}$.
Following recent progress in text-based approaches for the KGC task (yao2019kg, ; wang2022simkgc, ), we employ the strategy that individually encodes the textual descriptions of queries and answers using two BERT (devlin-etal-2019-bert, ) encoders, sharing pre-trained weights, and applying mean pooling:
$\mathbf{e}_{hr}=\text{Pool}(\text{BERT}_{\text{qry}}(\mathbf{x}_{hr})),~{}~{}\mathbf{e}_{t}=\text{Pool}(\text{BERT}_{\text{ans}}(\mathbf{x}_{t}))$ (15)
@@
$\{\pi_{hr,k},\mu_{hr,k},\sigma_{hr,k}\}_{k=1}^{K}=\text{MLP}(\mathbf{e}_{hr}),~{}~{}\{\pi_{t,k},\mu_{t,k},\sigma_{t,k}\}_{k=1}^{K}=\text{MLP}(\mathbf{e}_{t})$ (16)
It is important to note that the flexibility of our model allows for the use of various types of encoders, such as a multi-modal one, to further enhance expressiveness. We plan to explore these possibilities in future research.
We denote the overall encoder network with parameters $\phi$ as $h_{\phi}$. Integrating textual inputs not only enhances our model’s performance, but also provides deeper insights into the latent structure. This allows for the exploration of the mined communities through the descriptions of their constituent entities, the benefits of which will be demonstrated in the experiment section.